Identify the surface of the z2 = 4 + 4r2 by converting them into equations in the Cartesian form. Show the complete solutions.


1
Expert's answer
2022-06-01T02:44:06-0400
"z^2=4+4r^2=>z^2=4+4x^2+4y^2"

"-\dfrac{x^2}{1^2}-\dfrac{y^2}{1^2}+\dfrac{z^2}{2^2}=1"

The  hyperboloid of two sheets.